OpenStudy (anonymous):

2 over 5 (x - 2) = 4x. 2 over 9 9 negative 2 over 9 negative 9 over 2

OpenStudy (anonymous):

can someone tell me how to do the fraction part at least

OpenStudy (pixiedust1):

Okay! So basically, your asking\[\frac{ 2 }{ 5 }(x-2)=4x\] \[\frac{ 2 }{ 9 }\] \[9\] \[\frac{ -2 }{ 9 }\] \[\frac{ -9 }{ 2 }\]

OpenStudy (anonymous):

yes

OpenStudy (pixiedust1):

Okay! So, I'm gonna solve this the easy way :D Each of the answer options are the possible values of x, agree?

OpenStudy (anonymous):

i think so

OpenStudy (pixiedust1):

Okay! So, let's plug in each of the values of x for the x in the equation. Okay? So, \[\frac{ 2 }{ 5}(\frac{ 2 }{ 9}-2)=4*\frac{ 2 }{ 9 }\]

OpenStudy (pixiedust1):

So, we have to make both sides balanced. Both sides of the equal sign must match each other

OpenStudy (anonymous):

ok

OpenStudy (pixiedust1):

For this one, the left side equals \[-\frac{ 38}{ 45 }\] The right side equals \[\frac{ 8 }{ 9 }\] Are those two equal?

OpenStudy (anonymous):

no they are not

OpenStudy (pixiedust1):

Correct! Now, how do you think we would check the next one?

OpenStudy (anonymous):

do the same thing

OpenStudy (pixiedust1):

Great! Wanna give that a try for me? Walk me through how you think you would do that. If you get stuck, just ask for some help :D

OpenStudy (pixiedust1):

Did you want me to do one more before you try it yourself?

OpenStudy (anonymous):

ok when i did this problem and ai subtracted nine from 2 q=and got seven adn i multiplied 4 to nine and got 36 wht do i do with the 2 fifths

OpenStudy (anonymous):

is it the last one

OpenStudy (pixiedust1):

I'm sorry, I don't know what that means :( I think its a method that I haven't learned :( But you pretty much got the answer! You are correct! The answer is -2/9 Here's how we check out work. 2/5(-2/9-2)=-8/9 4*-2/9=-8/9 -8/9=-8/9

OpenStudy (pixiedust1):

Congratulations! So the third one is the answer!

OpenStudy (anonymous):

yes i hope

OpenStudy (pixiedust1):

What do you mean by "I hope" Is there something you are not understanding?

OpenStudy (anonymous):

no i got it now thank u

OpenStudy (pixiedust1):

Are you SURE? I ain't not believing you! :D